OpenStudy (anonymous):

Solve the problem. A tour group split into two groups when waiting in line for food at a fast food counter. The first group bought 8 slices of pizza and 5 soft drinks for $36.03. The second group bought 6 slices of pizza and 6 soft drinks for $31.32. How much does one slice of pizza cost? $1.91 per slice of pizza $3.31 per slice of pizza $2.41 per slice of pizza $2.81 per slice of pizza

OpenStudy (kropot72):

Let p be the cost of one slice of pizza and s be the cost of one soft drink. We can set up a pair of simultaneous equations to solve the question as follows: 8p + 5s = 36.03 .............(1) 6p + 6s = 31.32 .............(2) Can you now solve to find the value of p?

OpenStudy (kropot72):

Hint: Multiply both sides of equation (1) by 6 and both sides of equation (2) by 5. Then one equation can be subtracted from the other to eliminate the terms in s and give the value of p after simplification.

OpenStudy (wolf1728):

A) 8 p + 5 sd = 36.03 B) 6 p + 6 sd = 31.32 Multiply equation B by -8/6 B) -8 p - 8 sd = -41.76 Add that to Equation A and you'll be able to solve for the price of pizza and soft drinks.

OpenStudy (anonymous):

ah, $4.97?

OpenStudy (anonymous):

nope

OpenStudy (wolf1728):

4.97 isn't even one of the choices

OpenStudy (anonymous):

yeah, but that is what worked out, I'm back checking now

OpenStudy (kropot72):

48p + 30s = 216.18 ............(3) 30p + 30 s =156.60 ............(4) subtract (4) from (3) to eliminate the terms in s.

OpenStudy (anonymous):

3.31, found my mistake, thanks!

OpenStudy (wolf1728):

okay !!

OpenStudy (kropot72):

You are correct! Good work :)